If you invest 10,000 today in a hotel at 6.5% interest, compounded annually for 4 years. How much interest in interest did the hotel earn over this period of Time
1
Expert's answer
2020-10-14T18:18:02-0400

"I=P_0r"


"A=P0+I=P_0+P_0+P_0r=P_0(1+r)"


I is interest, A is principle plus interest, "P_0" is the principle, r is the rate

To get interest in four years we use simple interest formula;

"I=P_0\times r\times t(time)"

"I=10000\times 0.065\times 4"

"=2600"